# SENTRA: Selected-Next-Token Transformer for LLM Text Detection

### v1

@@ 1 Introduction
The problem of determining whether a text has been generated by an LLM or written by a human has been widely studied in both academia [Tang et al.(2024)Tang, Chuang, and Hu] and industry. Several commercial-level automated text detection systems have been developed, including GPTZero [Tian and Cui(2023)], Originality [Originality.AI(2025)], Sapling [Sapling AI(2025)], and Reality Defender [Reality Defender(2025)]. Although significant progress has been made in detecting LLM-generated text over the past several years, these systems remain far from perfect and are often unreliable. A major limitation is their brittleness: they can perform well on certain types of LLM-generated text but fail catastrophically in other cases [Dugan et al.(2024)Dugan, Hwang, Trhlík, Zhu, Ludan, Xu, Ippolito, and Callison-Burch]. This issue is particularly pronounced when operating in a real world scenario, where models must handle out-of-domain (OOD) data, different LLM generators, or various LLM "attacks" [Dugan et al.(2024)Dugan, Hwang, Trhlík, Zhu, Ludan, Xu, Ippolito, and Callison-Burch, Zhou et al.(2024)Zhou, He, and Sun]. Therefore, it is crucial to develop more generalizable methods that deliver reliable performance across these settings.
@@
The probability of a document under and LLM’s model can be measured by auto-regressively feeding the document’s tokens into the LLM and observing the predicted probabilities for each token. This process produces a sequence of values called selected-next-token-probabilities (SNTP) that has been extensively used in prior work on LLM-generated text detection [Guo et al.(2023)Guo, Zhang, Wang, Jiang, Nie, Ding, Yue, and Wu, Hans et al.(2024)Hans, Schwarzschild, Cherepanova, Kazemi, Saha, Goldblum, Geiping, and Goldstein, Verma et al.(2024)Verma, Fleisig, Tomlin, and Klein]. These prior works primarily rely on either heuristics (handcrafted functions) applied to SNTP sequences or linear models trained on expert-derived features [Hans et al.(2024)Hans, Schwarzschild, Cherepanova, Kazemi, Saha, Goldblum, Geiping, and Goldstein, Verma et al.(2024)Verma, Fleisig, Tomlin, and Klein]. In contrast, our proposed approach encodes SNTP sequences using a Transformer model pre-trained on unlabeled data, leveraging the expressivity of Transformers to directly learn a representation of the probability that a single or multiple LLMs assign to tokens in a document. More specifically, we propose SElected-Next-Token tRAnsformer (SENTRA), a method for detecting LLM-generated text that directly learns a detection function in a supervised manner from SNTP sequences. This method utilizes a novel Transformer-based architecture with a contrastive pre-training mechanism. The learned representation can be fine-tuned on labeled data to create a supervised model that distinguishes LLM-generated texts from human-written texts.
@@
For the LLM-text-detection task, supervised detectors have been shown to generalize poorly outside the training distribution [Dugan et al.(2024)Dugan, Hwang, Trhlík, Zhu, Ludan, Xu, Ippolito, and Callison-Burch]. Prior supervised methods typically leverage raw tokens as input and tend to overfit to token selections in a document. Heuristic or linear models on SNTP input have been shown to generalize well, but these simple models lack the expressivity to fully exploit the information in the SNTP sequences. Our SENTRA network addresses this issue by learning generalizable functions on SNTP. We show empirically that the supervised method presented in this paper generalizes to unseen domains better than both supervised and unsupervised baselines by leveraging our proposed Transformer-based architecture, thus demonstrating greater generalization to distribution shifts.
@@ 2 Related Work
With the rise of LLMs, significant research has been conducted on accurately detecting text generated by these models [Tang et al.(2024)Tang, Chuang, and Hu]. At a high level, these detectors can be categorized into three approaches: watermarking, unsupervised (or zero-shot) detection, and supervised detection. Watermarking generally relies on the LLM deliberately embedding identifiable traces in its output [Liu et al.(2025)Liu, Pan, Lu, Li, Hu, Zhang, Wen, King, Xiong, and Yu]. In this work, we focus on the general detection problem, including cases involving non-cooperative LLMs; therefore, we do not consider watermarking as a point of comparison. Unsupervised methods typically leverage metrics computed by an LLM on the target document. These methods can be further divided into white-box detection, where the candidate LLM is known [Mitchell et al.(2023)Mitchell, Lee, Khazatsky, Manning, and Finn], and black-box detection, where the candidate LLM is unknown [Tang et al.(2024)Tang, Chuang, and Hu]. Given our focus on the general detection problem, we prioritize black-box detection methods. Supervised methods, on the other hand, involve collecting a corpus of human-written and LLM-generated text samples, which are then used to train the detection models [Verma et al.(2024)Verma, Fleisig, Tomlin, and Klein, Soto et al.(2024)Soto, Koch, Khan, Chen, Bishop, and Andrews].
@@
Selected-next-token-probabilities (SNTP) have been widely used for LLM detection in both white and black box settings [Guo et al.(2023)Guo, Zhang, Wang, Jiang, Nie, Ding, Yue, and Wu, Hans et al.(2024)Hans, Schwarzschild, Cherepanova, Kazemi, Saha, Goldblum, Geiping, and Goldstein, Verma et al.(2024)Verma, Fleisig, Tomlin, and Klein]. Perplexity [Jelinek et al.(1977)Jelinek, Mercer, Bahl, and Baker] is a commonly used metric to evaluate an LLM’s ability to model a given dataset. In the context of AI detection, a lower perplexity score on a document indicates an LLM "fits" a document and this indicates a higher likelihood the document was LLM-generated. Conversely, a higher perplexity score suggests the LLM’s probability model does not fit or accurately represent the candidate text, implying a lower likelihood that the text was generated by the LLM [Guo et al.(2023)Guo, Zhang, Wang, Jiang, Nie, Ding, Yue, and Wu].
[…]
@@
Several publications have explored contrastive training for the LLM detection task [Bhattacharjee et al.(2023)Bhattacharjee, Kumarage, Moraffah, and Liu, Bhattacharjee et al.(2024)Bhattacharjee, Moraffah, Garland, and Liu, Soto et al.(2024)Soto, Koch, Khan, Chen, Bishop, and Andrews, Guo et al.(2024)Guo, He, Zhang, Zhang, Feng, Huang, and Ma]. These studies use contrastive pre-training for a text Transformer, which is chosen to be RoBERTa [Liu et al.(2019)Liu, Ott, Goyal, Du, Joshi, Chen, Levy, Lewis, Zettlemoyer, and Stoyanov] in many cases, to guide the network toward a representation more useful for LLM-generated text detection. Furthermore, many prior contrastive training strategies focus on identifying stylometric features [Soto et al.(2024)Soto, Koch, Khan, Chen, Bishop, and Andrews, Guo et al.(2024)Guo, He, Zhang, Zhang, Feng, Huang, and Ma], while other studies extract stylometric features directly and train classifiers using those features [Kumarage et al.(2023a)Kumarage, Garland, Bhattacharjee, Trapeznikov, Ruston, and Liu]. Rather than focusing on text representations, our method is mainly designed to produce useful SNTP representations and, thus, proposes a different contrastive pre-training scheme that compares textual representations with those of the SNTP Transformer.
@@
However, SNTP and supervised methods have been shown, both intuitively and empirically, to struggle with generalization to unseen domains [Li et al.(2024a)Li, Li, Cui, Bi, Wang, Wang, Yang, Shi, and Zhang, Roussinov et al.(2025)Roussinov, Sharoff, and Puchnina].
[…]
@@
Prior work has shown SNTP to be an effective input for identifying LLM generated text [Guo et al.(2023)Guo, Zhang, Wang, Jiang, Nie, Ding, Yue, and Wu, Hans et al.(2024)Hans, Schwarzschild, Cherepanova, Kazemi, Saha, Goldblum, Geiping, and Goldstein, Verma et al.(2024)Verma, Fleisig, Tomlin, and Klein], but they rely on relatively simple metrics or heuristics. In this paper, we propose a Transformer-based SENTRA model that learns a representation of SNTP sequences used for more effective training of detection models that better generalize to unseen domains.
@@ 3.1 Overview of the SENTRA Method
Prior work, reviewed in Section 2, has proposed various heuristic functions on these sequences that are useful in detecting LLM-generated text [Guo et al.(2023)Guo, Zhang, Wang, Jiang, Nie, Ding, Yue, and Wu, Hans et al.(2024)Hans, Schwarzschild, Cherepanova, Kazemi, Saha, Goldblum, Geiping, and Goldstein]. SENTRA replaces these heuristic functions on SNTP sequence(s) with a neural network, as shown in Figure 1 illustrating our proposed method. In particular, we leverage LLMs, each with parameters to produce SNTP sequences and for a candidate document with tokens using process in Equation 2. The sequences are concatenated to form input sequence . Note that in Figure 1, . In this work, we focus on the case. A allows the model to learn from similarities and deviations in SNTP sequences produced by LLMs. This comparison was a key idea in [Li et al.(2024b)Li, Wang, Cui, Bi, Shi, and Zhang], and following that work, we focus on the case where the two LLMs share a tokenizer. This allows the SNTP sequences to be aligned.
Instead of token embeddings often seen in Transformer architectures [Devlin et al.(2019)Devlin, Chang, Lee, and Toutanova], each token-indexed representation is independently projected using a fully connected layer.
where is the dense embedding representation, is the ReLU activation function, is the weight matrix, is the bias, and are learned positional embeddings. This transformation results in a representation of size for a single document. Note a learned [CLS] representation is pre-pended to the sequence before the positional embeddings are applied. This representation is passed through a bi-directional Transformer [Devlin et al.(2019)Devlin, Chang, Lee, and Toutanova] , as shown in Figure 1.
@@
In summary, SENTRA is the first Transformer-based encoder to systematically learn a useful representation of SNTP sequences. Similar to many Transformer-based approaches [Devlin et al.(2019)Devlin, Chang, Lee, and Toutanova, Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, Krueger, and Sutskever], that have traditionally used different modalities of input information, we demonstrate in Section 3.2 that our method can leverage large quantities of unlabeled data to enhance this learned representation.
@@ 3.2 SENTRA Contrastive Pre-Training
We further explore the pre-training of SENTRA using unlabeled text data and demonstrate in Section 4.4 that it significantly improves SENTRA’s performance. Notably, this pre-training scheme is reminiscent of CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, Krueger, and Sutskever]. Figure 2 illustrates our concept for pre-training SENTRA. We leverage off-the-shelf, pre-trained text representations to help SENTRA learn a useful representation of SNTP sequences. A document is encoded using both a pre-trained text encoder [Devlin et al.(2019)Devlin, Chang, Lee, and Toutanova, Liu et al.(2019)Liu, Ott, Goyal, Du, Joshi, Chen, Levy, Lewis, Zettlemoyer, and Stoyanov] and our SENTRA network, producing representations and . These representations are projected to a joint embedding space, and , using fully connected layers and for the text and SNTP representations respectively.
@@
In [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, Krueger, and Sutskever]’s work, the authors jointly trained text and image encoders from scratch. Unlike CLIP, which deals with text and images, we focus solely on text and on pre-training only the SENTRA SNTP encoder. To do this, we freeze a pre-trained text encoder and train only SENTRA and the contrastive embedding projections.
@@ 3.3 Implementation
We implement our SENTRA model with eight attention heads, eight layers, and a hidden dimension of for a total of 57M parameters. The Transformer architecture and positional embeddings follow the same definitions as in BERT [Devlin et al.(2019)Devlin, Chang, Lee, and Toutanova]. Before pre-training, the SENTRA parameters are randomly initialized. The frozen text encoder used for contrastive pre-training is initialized from RoBERTa [Liu et al.(2019)Liu, Ott, Goyal, Du, Joshi, Chen, Levy, Lewis, Zettlemoyer, and Stoyanov]. SENTRA is pre-trained on a 600K sample of Common Crawl data from RedPajama [Weber et al.(2024)Weber, Fu, Anthony, Oren, Adams, Alexandrov, Lyu, Nguyen, Yao, Adams, Athiwaratkun, Chalamala, Chen, Ryabinin, Dao, Liang, Ré, Rish, and Zhang]. Pre-training is conducted for epochs with a batch size of and a maximum token length of . We then continue contrastive training for epochs with a batch size of and a maximum token length of to pre-train the later position embeddings. The peak learning rate was set to for both phases. We use the AdamW [Loshchilov and Hutter(2019)] optimizer with a weight decay of and set the contrastive learning temperature to [Chen et al.(2020)Chen, Kornblith, Norouzi, and Hinton]. During fine-tuning, we initialize SENTRA from the pre-trained model, use a learning rate of , a weight decay of , and apply early stopping with a patience of two epochs on a validation dataset.
As shown in Figure 1, we implemented SENTRA with two SNTP sequences and therefore . Following Binoculars [Hans et al.(2024)Hans, Schwarzschild, Cherepanova, Kazemi, Saha, Goldblum, Geiping, and Goldstein], we use Falcon-7B and Falcon-7B-Instruct [Almazrouei et al.(2023)Almazrouei, Alobeidli, Alshamsi, Cappelli, Cojocaru, Debbah, Goffinet, Hesslow, Launay, Malartic, Mazzotta, Noune, Pannier, and Penedo] to produce these sequences. We used a sequence of two SNTP because Binoculars showed it is useful for the detector to compare both SNTP, and we used the Falcon models specifically because Binoculars showed they worked well [Hans et al.(2024)Hans, Schwarzschild, Cherepanova, Kazemi, Saha, Goldblum, Geiping, and Goldstein]. During SENTRA training, the SNTP sequences are precomputed and cached. At inference, the computational complexity is dominated by the Falcon models. Because we use the same LLMs as Binoculars [Hans et al.(2024)Hans, Schwarzschild, Cherepanova, Kazemi, Saha, Goldblum, Geiping, and Goldstein] and our SENTRA encoder is small, our method has the same order of complexity as Binoculars. See Appendix C for additional details.

### v2

@@ Abstract
SENTRA: Selected-Next-Token Transformer for LLM Text Detection
Mitchell Plyler Mozilla Corporation mlplyler@mozilla.com Yilun Zhang Mozilla Corporation tzhang@mozilla.com Alexander Tuzhilin New York University at2@stern.nyu.edu
Saoud Khalifah† Ciphero AI saoud@ciphero.ai Sen Tian† Ciphero AI sen@ciphero.ai
@@ 1 Introduction
The problem of determining whether a text has been generated by an LLM or written by a human has been widely studied in both academia Tang et al. (2024) and industry. Several commercial-level automated text detection systems have been developed, including GPTZero Tian and Cui (2023), Originality Originality.AI (2025), Sapling Sapling AI (2025), and Reality Defender Reality Defender (2025). Although significant progress has been made in detecting LLM-generated text over the past several years, these systems remain far from perfect and are often unreliable. A major limitation is their brittleness: they can perform well on certain types of LLM-generated text but fail catastrophically in other cases Dugan et al. (2024). This issue is particularly pronounced when operating in a real world scenario, where models must handle out-of-domain (OOD) data, different LLM generators, or various LLM "attacks" Dugan et al. (2024); Zhou et al. (2024). Therefore, it is crucial to develop more generalizable methods that deliver reliable performance across these settings.
@@
The probability of a document under an LLM’s model can be measured by auto-regressively feeding the document’s tokens into the LLM and observing the predicted probabilities for each token. This process produces a sequence of values called selected-next-token-probabilities (SNTP) that has been extensively used in prior work on LLM-generated text detection Guo et al. (2023); Hans et al. (2024); Verma et al. (2024). These prior works primarily rely on either heuristics (handcrafted functions) applied to SNTP sequences or linear models trained on expert-derived features Hans et al. (2024); Verma et al. (2024). In contrast, our proposed approach encodes SNTP sequences using a Transformer model pre-trained on unlabeled data, leveraging the expressivity of Transformers to directly learn a representation of the probability that a single or multiple LLMs assign to tokens in a document. More specifically, we propose SElected-Next-Token tRAnsformer (SENTRA), a method for detecting LLM-generated text that directly learns a detection function in a supervised manner from SNTP sequences. This method utilizes a novel Transformer-based architecture with a contrastive pre-training mechanism. The learned representation can be fine-tuned on labeled data to create a supervised model that distinguishes LLM-generated texts from human-written texts.
@@
For the LLM-text-detection task, supervised detectors have been shown to generalize poorly outside the training distribution Dugan et al. (2024). Prior supervised methods typically leverage raw tokens as input and tend to overfit to token selections in a document. Heuristic or linear models on SNTP input have been shown to generalize well, but these simple models lack the expressivity to fully exploit the information in the SNTP sequences. Our SENTRA network addresses this issue by learning generalizable functions on SNTP. We show empirically that the supervised method presented in this paper generalizes to unseen domains better than both supervised and unsupervised baselines by leveraging our proposed Transformer-based architecture, thus demonstrating greater generalization to distribution shifts.
@@ 2 Related Work
With the rise of LLMs, significant research has been conducted on accurately detecting text generated by these models Tang et al. (2024). At a high level, these detectors can be categorized into three approaches: watermarking, unsupervised (or zero-shot) detection, and supervised detection. Watermarking generally relies on the LLM deliberately embedding identifiable traces in its output Liu et al. (2025). In this work, we focus on the general detection problem, including cases involving non-cooperative LLMs; therefore, we do not consider watermarking as a point of comparison. Unsupervised methods typically leverage metrics computed by an LLM on the target document. These methods can be further divided into white-box detection, where the candidate LLM is known Mitchell et al. (2023), and black-box detection, where the candidate LLM is unknown Tang et al. (2024). Given our focus on the general detection problem, we prioritize black-box detection methods. Supervised methods, on the other hand, involve collecting a corpus of human-written and LLM-generated text samples, which are then used to train the detection models Verma et al. (2024); Soto et al. (2024).
@@
Selected-next-token-probabilities (SNTP) have been widely used for LLM detection in both white and black box settings Guo et al. (2023); Hans et al. (2024); Verma et al. (2024). Perplexity Jelinek et al. (1977) is a commonly used metric to evaluate an LLM’s ability to model a given dataset. In the context of AI detection, a lower perplexity score on a document indicates an LLM "fits" a document and this indicates a higher likelihood the document was LLM-generated. Conversely, a higher perplexity score suggests the LLM’s probability model does not fit or accurately represent the candidate text, implying a lower likelihood that the text was generated by the LLM Guo et al. (2023).
Some detectors use multiple sequences of SNTP for the detection task Verma et al. (2024); Hans et al. (2024). Verma et al. (2024) leveraged SNTPs from two Markov models, along with an LLM’s SNTP, extracted features, and a forward feature selection scheme as inputs to a linear classifier. In contrast to Guo et al. (2023), Hans et al. (2024) argued that relying solely on the perplexity score for LLM-generated content detection can be misleading. Although human-authored text generally results in higher perplexity, prompts can significantly influence perplexity values. The authors highlighted the "capybara problem", where the absence of a prompt can cause an LLM-generated response to have higher perplexity, leading to false detections. They addressed this issue by introducing cross-perplexity as a normalizing factor to calibrate for prompts that yield high perplexity.
DetectGPT is an unsupervised method based on the idea that texts generated by LLMs tend to "occupy negative curvature regions of the model’s log probability function" Mitchell et al. (2023). The method generates perturbations of the sample text using a smaller model and compares the log probability of the sample text to that of the perturbations. Fast-DetectGPT replaces the perturbations in DetectGPT with a more efficient sampling step Bao et al. (2024). Nguyen-Son et al. (2024) observed that the similarity between a sample and its counterpart generation is notably higher than the similarity between the counterpart and another independent regeneration. They demonstrated that this difference in similarity is useful for detection.
The most common supervised baseline for LLM-generated text detection is a RoBERTa classifier Liu et al. (2019) trained on a corpus of labeled text, where each document is marked as either human-written or LLM-generated. Several studies have expanded on this approach to supervised text-based classification. Yu et al. (2024) trained a feed-forward classifier with two hidden layers using intrinsic features derived from Transformer hidden states, determined via KL-divergence. Tian et al. (2024) address the challenge of detecting short texts by treating short samples in the training corpus as partially "unlabeled". Hu et al. (2023) employed adversarial learning to improve the robustness of their RoBERTa-based classifier against paraphrase attacks.
@@
Several publications have explored contrastive training for the LLM detection task Bhattacharjee et al. (2023, 2024); Soto et al. (2024); Guo et al. (2024). These studies use contrastive pre-training for a text Transformer, which is chosen to be RoBERTa Liu et al. (2019) in many cases, to guide the network toward a representation more useful for LLM-generated text detection. Furthermore, many prior contrastive training strategies focus on identifying stylometric features Soto et al. (2024); Guo et al. (2024), while other studies extract stylometric features directly and train classifiers using those features Kumarage et al. (2023a). Rather than focusing on text representations, our method is mainly designed to produce useful SNTP representations and, thus, proposes a different contrastive pre-training scheme that compares textual representations with those of the SNTP Transformer.
@@
However, SNTP and supervised methods have been shown, both intuitively and empirically, to struggle with generalization to unseen domains Li et al. (2024a); Roussinov et al. (2025).
For instance, Lai et al. (2024) applied adaptive ensemble algorithms to enhance model performance in OOD scenario. Meanwhile, Guo et al. (2024) and Soto et al. (2024), recognizing the limited number of widely adopted general-purpose AI assistants, proposed to train an embedding model to learn the writing style of LLMs, and thereby improving the detection accuracy.
@@
Prior work has shown SNTP to be an effective input for identifying LLM generated text Guo et al. (2023); Hans et al. (2024); Verma et al. (2024), but they rely on relatively simple metrics or heuristics. In this paper, we propose a Transformer-based SENTRA model that learns a representation of SNTP sequences used for more effective training of detection models that better generalize to unseen domains.
@@ 3.1 Overview of the SENTRA Method
Prior work, reviewed in Section 2, has proposed various heuristic functions on these sequences that are useful in detecting LLM-generated text Guo et al. (2023); Hans et al. (2024). SENTRA replaces these heuristic functions on SNTP sequence(s) with a neural network, as shown in Figure 1 illustrating our proposed method. In particular, we leverage LLMs, each with parameters to produce SNTP sequences and for a candidate document with tokens using process in Equation 2. The sequences are concatenated to form input sequence . Note that in Figure 1, . In this work, we focus on the case. Setting allows the model to learn from similarities and deviations in SNTP sequences produced by LLMs. This comparison was a key idea in Li et al. (2024b), and following that work, we focus on the case where the two LLMs share a tokenizer. This allows the SNTP sequences to be aligned.
Instead of token embeddings often seen in Transformer architectures Devlin et al. (2019), each token-indexed representation is independently projected using a fully connected layer.
where is the dense embedding representation, is the ReLU activation function, is the weight matrix, is the bias, and are learned positional embeddings. This transformation results in a representation of size for a single document. Note a learned [CLS] representation is pre-pended to the sequence before the positional embeddings are applied. This representation is passed through a bi-directional Transformer Devlin et al. (2019) , as shown in Figure 1.
@@
In summary, SENTRA is the first Transformer-based encoder to systematically learn a useful representation of SNTP sequences. Similar to many Transformer-based approaches Devlin et al. (2019); Radford et al. (2021), that have traditionally used different modalities of input information, we demonstrate in Section 3.2 that our method can leverage large quantities of unlabeled data to enhance this learned representation.
@@ 3.2 SENTRA Contrastive Pre-Training
We further explore the pre-training of SENTRA using unlabeled text data and demonstrate in Section 4.4 that it significantly improves SENTRA’s performance. Notably, this pre-training scheme is reminiscent of CLIP Radford et al. (2021). Figure 2 illustrates our concept for pre-training SENTRA. We leverage off-the-shelf, pre-trained text representations to help SENTRA learn a useful representation of SNTP sequences. A document is encoded using both a pre-trained text encoder Devlin et al. (2019); Liu et al. (2019) and our SENTRA network, producing representations and . These representations are projected to a joint embedding space, and , using fully connected layers and for the text and SNTP representations respectively.
@@
In Radford et al. (2021)’s work, the authors jointly trained text and image encoders from scratch. Unlike CLIP, which deals with text and images, we focus solely on text and on pre-training only the SENTRA SNTP encoder. To do this, we freeze a pre-trained text encoder and train only SENTRA and the contrastive embedding projections.
@@ 3.3 Implementation
We implement our SENTRA model with eight attention heads, eight layers, and a hidden dimension of for a total of 57M parameters. The Transformer architecture and positional embeddings follow the same definitions as in BERT Devlin et al. (2019). Before pre-training, the SENTRA parameters are randomly initialized. The frozen text encoder used for contrastive pre-training is initialized from RoBERTa Liu et al. (2019). SENTRA is pre-trained on a 600K sample of Common Crawl data from RedPajama Weber et al. (2024). Pre-training is conducted for epochs with a batch size of and a maximum token length of . We then continue contrastive training for epochs with a batch size of and a maximum token length of to pre-train the later position embeddings. The peak learning rate was set to for both phases. We use the AdamW Loshchilov and Hutter (2019) optimizer with a weight decay of and set the contrastive learning temperature to Chen et al. (2020). During fine-tuning, we initialize SENTRA from the pre-trained model, use a learning rate of , a weight decay of , and apply early stopping with a patience of two epochs on a validation dataset.
As shown in Figure 1, we implemented SENTRA with two SNTP sequences and therefore . Following Binoculars Hans et al. (2024), we use Falcon-7B and Falcon-7B-Instruct Almazrouei et al. (2023) to produce these sequences. We used a sequence of two SNTP because Binoculars showed it is useful for the detector to compare both SNTP, and we used the Falcon models specifically because Binoculars showed they worked well Hans et al. (2024). During SENTRA training, the SNTP sequences are precomputed and cached. At inference, the computational complexity is dominated by the Falcon models. Because we use the same LLMs as Binoculars Hans et al. (2024) and our SENTRA encoder is small, our method has the same order of complexity as Binoculars. See Appendix C for additional details.